# TRACE: Learning to Compute on Graphs

### v1

@@ Abstract
Learning to compute—the ability to model the functional behavior of a computational graph—is a fundamental challenge for graph representation learning. Yet, the dominant paradigm is architecturally mismatched for this task. This flawed assumption, central to mainstream message passing neural networks (MPNNs) and their conventional Transformer-based counterparts, prevents models from capturing the position-aware, hierarchical nature of computation. To resolve this, we introduce TRACE, a new paradigm built on an architecturally sound backbone and a principled learning objective. First, TRACE employs a Hierarchical Transformer that mirrors the step-by-step flow of computation, providing a faithful architectural backbone that replaces the flawed permutation-invariant aggregation. Second, we introduce function shift learning, a novel objective that decouples the learning problem. Instead of predicting the complex global function directly, our model is trained to predict only the function shift—the discrepancy between the true global function and a simple local approximation that assumes input independence. We validate this paradigm on electronic circuits, one of the most complex and economically critical classes of computational graphs. Across a comprehensive suite of benchmarks, TRACE substantially outperforms all prior architectures. These results demonstrate that our architecturally-aligned backbone and decoupled learning objective form a more robust paradigm for the fundamental challenge of learning to compute on graphs.
00footnotetext: Equal contribution.
1 Introduction
@@
Computational graphs provide a fundamental abstraction for modeling computation. As directed graphs of nodes representing operations and variables, they capture the flow of computation and are crucial in domains ranging from control data flow graphs (CDFGs) in software engineering to electronic circuits in hardware design. Accurately modeling the computational behavior of these graphs is therefore a critical enabler for high-impact applications, including performance prediction (Chen et al., 2024; Xie et al., 2022; Zhang et al., 2020; Mendis et al., 2019), verification (Li et al., 2023; Selsam et al., 2018; Zhang et al., 2021), and optimization (Zuo et al., 2023; Mirhoseini et al., 2021).
@@
This need for functional modeling has driven recent work in graph representation learning, with approaches largely divided into two families: message passing neural networks (MPNNs) and Graph Transformers. These models have been increasingly applied to capture the functionality of diverse computational graph modalities, with a significant body of work focusing on the particularly challenging domain of hardware design, including Register Transfer Level (RTL) graphs (Fang et al., 2025c; b), And-Inverter Graphs (AIGs) (Shi et al., 2023; 2024; Zheng et al., 2025; Khan et al., 2025; Wu et al., 2025; Wang et al., 2024; Liu et al., 2024), and post-mapping (PM) netlists (Shi et al., 2025b). MPNNs provide a general framework (Gilmer et al., 2017) based on a paradigm of aggregating and updating node features:
@@
$\mathbf{x}_{i}^{\prime}=\gamma\left(\mathbf{x}_{i},\,\underset{j\in\mathcal{N}(i)}{\square}\,\varphi\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{e}_{j,i}\right)\right),$ (1)
@@
where $\square$ denotes a permutation-invariant aggregator (e.g., sum, mean, max) and the update function $\gamma$ and message function $\varphi$ are differentiable functions. In contrast, Transformer-based methods either flatten the graph into a sequence for global self-attention (Fang et al., 2025c; b), or incorporate graph structure via attention masks (Shi et al., 2024; Zheng et al., 2025; Fang et al., 2025a).
@@
Figure 1: The architectural failure of MPNNs on computational graphs. Left: The permutation-invariant aggregation in MPNNs cannot distinguish between ordered inputs (e.g., A,B vs B,A), yielding the same incorrect embedding for a position-aware operator like MUX. Right: Our approach processes inputs as an ordered sequence, enabling position-awareness and capturing operator-specific interactions.
@@
Despite their success on general graph domains, we argue that these architectures are fundamentally ill-suited for computation due to a fundamental mismatch between their properties and the nature of computational graphs (as shown in Fig. 1):
  1. 1.
@@
  2. 2.
    Vanilla Transformers are Hierarchy-Agnostic: Fully-connected Transformers flatten the graph into a sequence, destroying the explicit hierarchy and connectivity crucial for modeling computation. For an expression like $y=f(g(a,b),c)$, this architecture fails to capture the intermediate dependency on $g(a,b)$ and cannot guarantee a correct computational trace.
  3. 3.
@@
These architectural flaws highlight a failure to model the step-by-step flow of computation. Beyond this, however, a deeper challenge lies in capturing global function that emerges from the graph’s overall topology. Even with simple components (e.g., AND and NOT gates in AIGs), a graph’s overall functionality can become highly complex due to reconvergent dependencies. For instance, consider $c=a\land b$ with $a=x\land y$ and $b=y\land z$, where $x,y,z\sim\mathcal{B}(p)$ (a Bernoulli distribution with parameter $p$). Locally, $a,b\sim\mathcal{B}(p^{2})$, and ignoring reconvergence, one would predict $c\sim\mathcal{B}(p^{4})$. However, since $y$ appears in both $a$ and $b$, $a$ and $b$ are correlated, shifting the true distribution to $c\sim\mathcal{B}(p^{3})$. This function shift—from $\mathcal{B}(p^{4})$ to $\mathcal{B}(p^{3})$—demonstrates how dependencies fundamentally alter functional behavior. Previous works that supervise directly on a node’s final global function implicitly bundle this effect into the embeddings, making it difficult for the model to distinguish true functional dependencies from spurious correlations.
@@
Our work addresses these challenges with TRACE, a Transformer for Reasoning about Algebraic and Computational Expressions, which presents a two-fold solution. First, to resolve the local architectural flaws of prior models, TRACE employs a hierarchical Transformer. Inspired by prefix notation, we represent each computation step as an ordered sequence, [operator, input_1, input_2, …], which is processed by a Transformer encoder with positional encoding (Figure 1). By applying this process recursively according to the graph’s logical dependencies, TRACE learns a faithful, position-aware representation of each computational step. Second, to capture the global function, we introduce function shift learning, a novel objective that explicitly models the discrepancy between local and global functions. This allows the model to disentangle a node’s intrinsic behavior from the contextual effects imposed by the wider graph topology.
@@
We demonstrate the effectiveness of TRACE on electronic circuits—a particularly challenging and representative class of computational graphs. Our experiments span a broad spectrum of circuit modalities (RTL, AIG, and PM netlists) and tasks (contrastive and predictive) across several standard benchmarks, including ITC (Corno et al., 2002), OpenCores (Albrecht, 2005), ISCAS ’89 (Brglez et al., 1989), ForgeEDA (Shi et al., 2025a), and DeepCircuitX (Li et al., 2025). The results are unequivocal: TRACE consistently and substantially outperforms prior approaches across all settings. This establishes TRACE not only as a new state of the art for circuit analysis, but as a more robust and architecturally sound paradigm for learning on computational graphs.
2 Background
@@
Figure 2: Illustration of a Computational Graph and Its Computation Process. This figure demonstrates the dual role of nodes within a computational graph. Step 1 shows nodes 1 and 2 functioning as operators to compute the expressions $S\land B$ and $\lnot S$, respectively. As the computation progresses to Step 2, these nodes transition to representing the intermediate variables that hold the results of these operations in step 1, which are then passed to subsequent operators for further computation.
2.1 Computational Graphs
@@
In our work, computational graphs are defined as directed graphs where nodes represent either input variables or operators, and edges signify the flow of data. As depicted in Figure 2, this structure allows a single node to serve a dual role: it can be an operator receiving data from its source nodes and an intermediate variable whose output is consumed by other operators. Specifically, for any directed edge from a source to a target node, the target always represents an operator, while the source represents a variable—either an initial input variable or a temporary intermediate variable resulting from a previous operation. This representation effectively models the data dependencies and computational flow within the system. We focus on three types of computational graphs from the front-end of the electronic design automation (EDA) flow: Register-Transfer Level (RTL) graphs, And-Inverter Graphs (AIGs), and Post-Mapping (PM) netlists. The primary distinction among these graph types lies in their operators. AIGs are composed of a minimal set of basic operators (AND and NOT gates), while RTL and PM netlists feature more complex, higher-level operators, such as multiplexers where $\texttt{MUX(S,A,B)}=(S\land B)\lor(A\land\lnot S)$. For a more detailed description of the computational graphs used in this paper, please refer to Appendix B.
A notable property of computational graphs, and one that differentiates them from general graphs, is the distribution of node in-degrees. Unlike the often long-tailed distribution found in general graphs, the in-degree of a node in a computational graph is directly related to its operator type, leading to a stable and predictable distribution. This inherent property, detailed further in Appendix D, contributes to the substantially reduced padding overhead of our proposed model.
2.2 Message Passing Neural Networks
@@
Message Passing Neural Networks (MPNNs) are a dominant architectural paradigm for function learning on circuit graphs. These models can be broadly categorized into two types: synchronous and asynchronous. Synchronous MPNNs (Wu et al., 2023a; Liu et al., 2024; Wu et al., 2025; Deng et al., 2024) process all message-passing updates in parallel, a strategy designed for computational efficiency. In contrast, asynchronous MPNNs (Li et al., 2022; Shi et al., 2023; 2025b; Khan et al., 2025; Wang et al., 2024) mimic the logic simulation process by updating node representations sequentially, following a topological order. This approach aims to capture effective functional representations by emulating the data flow. However, as illustrated in Figure 1, both synchronous and asynchronous MPNNs rely on the conventional message passing paradigm (Gilmer et al., 2017), which fundamentally struggles to capture the operator-specific and position-aware interactions among input variables.
2.3 Graph Transformers
Despite the widespread use of MPNNs, they suffer from inherent limitations, including difficulty in capturing long-range dependencies and susceptibility to issues like over-smoothing (Li et al., 2018) and over-squashing (Alon & Yahav, 2020). These drawbacks have motivated a shift towards Graph Transformers, which leverage global attention mechanisms to address these limitations. Existing graph transformer models can be classified into two main categories: fully-connected transformers and edge-masked transformers.
Fully-connected transformers, such as Rampášek et al. (2022); Wu et al. (2023c; b) for general graphs and Fang et al. (2025c; b) for circuit graphs, treat the graph as a flattened sequence of nodes. While this enables global self-attention, it inadvertently leads to a loss of the critical hierarchical structure and intermediate dependencies that are inherent to computational graphs. In contrast, edge-masked transformers, including DeepGate3 (Shi et al., 2024), DeepGate4 (Zheng et al., 2025), and NetTAG (Fang et al., 2025a), integrate the graph’s topology by using the adjacency matrix to mask the attention mechanism. However, this approach often reduces the model to a Graph Attention Network (GAT)-like attention mechanism, inheriting the limitations of traditional MPNNs.
3 Method
3.1 Overview
Our framework learns circuit functionality using a Hierarchical Transformer backbone trained with two objectives: a predictive task and a contrastive task. The Hierarchical Transformer, detailed in Section 3.2, provides an architectural backbone that mirrors the circuit’s computational flow, replacing the conventional message-passing paradigm. For the predictive task, we introduce Function Shift Learning (FSL) in Section 3.3, a novel objective that explicitly models the discrepancy between a circuit’s local and global functions.
@@ 3.1 Overview
Figure 3: Overview of our proposed framework. Left: A circuit graph, represented in both a graph view and its equivalent prefix notation, is encoded by a Hierarchical Transformer to model the computation process. Right: For predictive tasks, we introduce Function Shift Learning (FSL). Instead of directly regressing the global function, the model captures the difference between the global and local functions: $y^{FSL}=y^{global}-y^{local}$. For contrastive tasks, we pull embeddings of equivalent circuits closer while pushing apart those of functionally different circuits.
3.2 Hierarchical Transformer
@@
To address the limitations of MPNNs and Graph Transformers discussed in Section 1 and  2, we propose a new paradigm for encoding the computational graph, with a position-aware Hierarchical Transformer that enables the operator-specific interaction among input variables, faithfully mirroring the computational process illustrated in Figure 2.
@@
A computational graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$, consists of primary input (PI) nodes, which have an in-degree of zero, and operator nodes, whose in-degree is determined by their type. We begin by computing the logic level of each node in topological order [^1] To handle cyclic graphs, such as sequential AIGs, we follow Khan et al. (2025) by treating all flip-flops or registers as pseudo Primary Inputs (PIs) and removing the feedback edges to compute the logical level. as follows:
$level(v)=\begin{cases}0&\text{if $v$ is PI}\\ 1+\max\limits_{(u,v)\in\mathbf{E}}level(u)&\text{otherwise}\end{cases}$ (2)
@@
The logic level defines the computational dataflow through the graph. Inspired by asynchronous MPNNs, we process nodes level by level. First, Primary Input (PI) nodes (level 0) are initialized with their input distribution, e.g., $\mathcal{B}(p)$, a Bernoulli distribution with parameter $p$. Then, for the nodes $v_{0}^{k},v_{1}^{k},\dots,v_{n_{k}}^{k}$ at subsequent level $k>0$, we gather their respective sets of direct predecessors, $\mathcal{N}(v_{0}^{k}),\dots,\mathcal{N}(v_{n_{k}}^{k})$, where each set is defined as $\mathcal{N}(v_{i}^{k})=\{u_{j}\in\mathbf{V}\mid(u_{j},v_{i}^{k})\in\mathbf{E}\}$.
@@
Inspired by prefix notation, we represent each computation at level $k$ as an ordered sequence, $\mathbf{v}_{i}^{k}$. This sequence is constructed by placing the operator $v_{i}^{k}$ at the head, followed by its ordered input nodes from $\mathcal{N}(v_{i}^{k})$:
$\mathbf{v}_{i}^{k}=[v_{i}^{k},u_{j_{1}},u_{j_{2}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}],\text{ where }u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}\in\mathcal{N}(v_{i}^{k})$ (3)
@@
$v_{i}^{k}=\text{Transformer}(\mathbf{v}_{i}^{k}+\mathbf{pos})[0],$ (4)
@@
This paradigm offers several key advantages. First, positional encodings enable position-aware aggregation, a critical feature for order-dependent operations, which contrasts with the permutation-invariant nature of standard message-passing schemes. Second, the self-attention mechanism facilitates rich interactions among all source nodes $u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}$], faithfully mirroring the computational dataflow of an operator, as illustrated in Figure 2. Finally, the sequence length for the Transformer depends only on a node’s in-degree. Unlike general graphs, which often have a long-tailed degree distributions that incur significant overhead, circuit graphs have small, tightly bounded in-degrees. This inherent property, detailed further in Appendix D, contributes to the substantially reduced padding overhead of our proposed model, as we discussed in Section 2 and Appendix D.
An Alternative View: Hierarchical Transformer on Prefix Notation
Our proposed method can also be interpreted as a Hierarchical Transformer operating on prefix notation, in a way that follows the logical dependencies inherent in the computational graph. This perspective highlights how the model processes the graph by mirroring the step-by-step evaluation of an expression. A computational graph can be converted into a prefix notation string through a pre-order traversal on its reversed edges. As illustrated in Figure 3, our proposed Hierarchical Transformer, unlike methods that simply flatten the graph into a single sequence, inherently preserves the nested, hierarchical structure of dependencies within each computation step, which allows it to align naturally with the actual flow of computation. This view also underscores the generalizability of our approach to other sequence-based problems, such as hardware model-checking, where the problem is often represented in Btor2 format (ArminBiere et al., 2018), a prefix-style representation of bit-vector formulas.
3.3 Function Shift Learning on Predictive Task
Logic-1 probability prediction is a task widely studied in prior works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2024; Zheng et al., 2025; Shi et al., 2025b; Liu et al., 2024), as it serves as a key indicator of a model’s ability to capture circuit functionality. The logic-1 probability corresponds to the global function of a circuit, as defined in Definition 1. This function can be highly complex due to reconvergent dependencies (see Section 1), and computing it directly requires enumerating the joint distribution of all inputs, which incurs an exponential cost of $O(2^{k})$.

### v2

# TRACE: Learning to Compute on Circuit Graphs

@@ Abstract.
Learning to compute, the ability to model the functional behavior of a circuit graph, is a fundamental challenge for graph representation learning. Yet, the dominant paradigm is architecturally mismatched for this task. This flawed assumption, central to mainstream message passing neural networks (MPNNs) and their conventional Transformer-based counterparts, prevents models from capturing the position-aware, hierarchical nature of computation. To resolve this, we introduce TRACE, a new paradigm built on an architecturally sound backbone and a principled learning objective. First, TRACE employs a Hierarchical Transformer that mirrors the step-by-step flow of computation, providing a faithful architectural backbone that replaces the flawed permutation-invariant aggregation. Second, we introduce function shift learning, a novel objective that decouples the learning problem. Instead of predicting the complex global function directly, our model is trained to predict only the function shift, the discrepancy between the true global function and a simple local approximation that assumes input independence. We validate this paradigm on various circuits modalities, including Register Transfer Level graphs, And-Inverter Graphs and post-mapping netlists. Across a comprehensive suite of benchmarks, TRACE substantially outperforms all prior architectures. These results demonstrate that our architecturally-aligned backbone and decoupled learning objective form a more robust paradigm for the fundamental challenge of learning the functional behavior of a circuit graph.
conference: ; ; 00footnotetext: Equal contribution.
1. Introduction
@@
Circuit graphs provide a fundamental abstraction for modeling computation. As directed graphs of nodes representing operations and variables, they capture the flow of computation and are crucial in various circuit modalities. Accurately modeling the computational behavior of circuit graphs is therefore a critical enabler for high-impact applications, including performance prediction (Chen et al., 2024; Xie et al., 2022; Zhang et al., 2020; Mendis et al., 2019), verification (Li et al., 2023; Selsam et al., 2018; Zhang et al., 2021), and optimization (Zuo et al., 2023; Mirhoseini et al., 2021).
@@
This need for functional modeling has driven recent work in graph representation learning, with approaches largely divided into two families: message passing neural networks (MPNNs) and Graph Transformers. These models have been increasingly applied to capture the functionality of diverse computational graph modalities, with a significant body of work focusing on the particularly challenging domain of hardware design, including Register Transfer Level (RTL) graphs (Fang et al., 2025c, b), And-Inverter Graphs (AIGs) (Shi et al., 2023, 2024; Zheng et al., 2025; Khan et al., 2025; Wu et al., 2025; Wang et al., 2024; Liu et al., 2024), and post-mapping (PM) netlists (Shi et al., 2025b). MPNNs provide a general framework (Gilmer et al., 2017) based on a paradigm of aggregating and updating node features:
@@
(1) $\mathbf{x}_{i}^{\prime}=\gamma\left(\mathbf{x}_{i},\,\underset{j\in\mathcal{N}(i)}{\square}\,\varphi\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{e}_{j,i}\right)\right),$
@@
where $\square$ denotes a permutation-invariant aggregator (e.g., sum, mean, max) and the update function $\gamma$ and message function $\varphi$ are differentiable functions. In contrast, Transformer-based methods either flatten the graph into a sequence for global self-attention (Fang et al., 2025c, b), or incorporate graph structure via attention masks (Shi et al., 2024; Zheng et al., 2025; Fang et al., 2025a).
@@
Figure 1. The architectural failure of MPNNs on circuit graphs. Top: The permutation-invariant aggregation in MPNNs cannot distinguish between ordered inputs (e.g., A,B vs B,A), yielding the same incorrect embedding for a position-aware operator like MUX. Bottom: Our approach processes inputs as an ordered sequence, enabling position-awareness and capturing operator-specific interactions.
@@
Despite their success on general graph domains, we argue that these architectures are ill-suited for computation due to a fundamental mismatch between their properties and the nature of computation process in circuit graphs (as shown in Fig. 1):
  1. (1)
@@
  2. (2)
    Vanilla Transformers are Hierarchy-Agnostic: Fully-connected Transformers flatten the graph into a sequence, destroying the explicit hierarchy and connectivity for modeling computation. For an expression like $y=f(g(a,b),c)$, this architecture fails to capture the intermediate dependency on $g(a,b)$ and cannot guarantee a correct computational trace.
  3. (3)
@@
These architectural flaws highlight a failure to model the step-by-step flow of computation. Beyond this, however, a deeper challenge lies in capturing global function that emerges from the graph’s overall topology. Even with simple components (e.g., AND and NOT gates in AIGs), a graph’s overall functionality can become highly complex due to reconvergent dependencies. For instance, consider $c=a\land b$ with $a=x\land y$ and $b=y\land z$, where $x,y,z\sim\mathcal{B}(p)$ (a Bernoulli distribution with parameter $p$). Locally, $a,b\sim\mathcal{B}(p^{2})$, and ignoring reconvergence, one would predict $c\sim\mathcal{B}(p^{4})$. However, since $y$ appears in both $a$ and $b$, $a$ and $b$ are correlated, shifting the true distribution to $c\sim\mathcal{B}(p^{3})$. This function shift, from $\mathcal{B}(p^{4})$ to $\mathcal{B}(p^{3})$, demonstrates how dependencies fundamentally alter functional behavior. Previous works that supervise directly on a node’s final global function implicitly bundle this effect into the embeddings, making it difficult for the model to distinguish true functional dependencies from spurious correlations.
@@
Our work addresses these challenges with TRACE, a Transformer for Reasoning about Algebraic and Computational Expressions, which presents a two-fold solution. First, to resolve local architectural flaws of prior models, TRACE employs a hierarchical Transformer. Inspired by prefix notation, we represent a computation step as an ordered sequence, [operator, input_1, input_2, …], processed by a Transformer encoder with positional encoding (Figure 1). Applying this process recursively according to graph’s logical dependencies, TRACE learns a faithful, position-aware representation of each computational step. Second, to capture the global function, we introduce function shift learning, a novel objective that explicitly models the discrepancy between local and global functions, allowing the model to disentangle a node’s intrinsic behavior from the contextual effects imposed by wider graph topology.
@@
Our experiments span a broad spectrum of circuit modalities (RTL, AIG, and PM netlists) and tasks (contrastive and predictive) across several standard benchmarks, including ITC (Corno et al., 2002), OpenCores (Albrecht, 2005), ISCAS ’89 (Brglez et al., 1989), ForgeEDA (Shi et al., 2025a), and DeepCircuitX (Li et al., 2025). The results are unequivocal: TRACE consistently and substantially outperforms prior approaches across all settings. This establishes TRACE not only as a new state of the art for circuit analysis, but as a more robust and architecturally sound paradigm for learning on circuit graphs.
2. Background
@@
Figure 2. Illustration of a circuit graph and its computation process. This figure demonstrates the dual role of nodes within a circuit graph. Step 1 shows nodes 1 and 2 functioning as operators to compute the expressions $S\land B$ and $\lnot S$, respectively. As computation progresses to Step 2, these nodes transition to representing the intermediate variables that hold the results of these operations in step 1, which are then passed to subsequent operators for further computation.
2.1. Circuit Graphs
@@
In our work, circuit graphs are defined as directed graphs where nodes represent either input variables or operators, and edges signify data flow. As depicted in Figure 2, this structure allows a single node to serve dual roles: it can be an operator receiving data from its source nodes and an intermediate variable whose output is consumed by other operators. Specifically, for any directed edge from a source to a target node, the target always represents an operator, while the source represents a variable—either an initial input variable or a temporary intermediate variable resulting from a previous operation. This representation effectively models the data dependencies and computational flow within the system. We focus on three types of circuit graphs from the front-end of the electronic design automation (EDA) flow: Register-Transfer Level (RTL) graphs, And-Inverter Graphs (AIGs), and Post-Mapping (PM) netlists.
2.2. Message Passing Neural Networks
@@
Message Passing Neural Networks (MPNNs) are a dominant architectural paradigm for function learning on circuit graphs. These models can be broadly categorized into two types: synchronous and asynchronous. Synchronous MPNNs (Wu et al., 2023a; Liu et al., 2024; Wu et al., 2025; Deng et al., 2024) process all message-passing updates in parallel, a strategy designed for computational efficiency. In contrast, asynchronous MPNNs (Li et al., 2022; Shi et al., 2023, 2025b; Khan et al., 2025; Wang et al., 2024) mimic the logic simulation process by updating node representations sequentially, following a topological order, which aims to capture effective functional representations by emulating the data flow. Such asynchronous processing may introduces a potential limitation for processing circuits with significant logic depth. This limitation is often mitigated in practice, as real-world timing constraints typically restrict the maximum number of logic levels. Despite their architectural differences, both synchronous and asynchronous MPNNs rely on the conventional message passing paradigm (Gilmer et al., 2017) (Figure 1). This paradigm fundamentally struggles to capture the operator-specific and position-aware interactions among input variables.
@@
Figure 3. Overview of our proposed framework. Left: A circuit graph, represented in both a graph view and its equivalent prefix notation, is encoded by a Hierarchical Transformer to model the computation process. Right: For predictive tasks, we introduce Function Shift Learning (FSL). Instead of directly regressing the global function, the model captures the difference between the global and local functions: $y^{FSL}=y^{global}-y^{local}$.
2.3. Graph Transformers
Despite the widespread use of MPNNs, they suffer from inherent limitations, including difficulty in capturing long-range dependencies and susceptibility to over-smoothing (Li et al., 2018) and over-squashing (Alon and Yahav, 2020). These drawbacks have motivated a shift towards Graph Transformers, which leverage global attention to address them. Existing graph transformers fall into two categories: fully-connected transformers and edge-masked transformers.
Fully-connected transformers, such as Rampášek et al. (2022); Wu et al. (2023c, b) for general graphs and Fang et al. (2025c, b) for circuit graphs, treat the graph as a flattened sequence of nodes. While this enables global self-attention, it inadvertently leads to a loss of the critical hierarchical structure and intermediate dependencies that are inherent to circuit graphs. In contrast, edge-masked transformers, including DeepGate3 (Shi et al., 2024), DeepGate4 (Zheng et al., 2025), and NetTAG (Fang et al., 2025a), integrate the graph’s topology by using the adjacency matrix to mask the attention mechanism. However, this approach often reduces the model to a Graph Attention Network (GAT)-like attention mechanism, inheriting the limitations of traditional MPNNs.
3. Method
3.1. Hierarchical Transformer
@@
To address the limitations of MPNNs and Graph Transformers discussed in Section 1 and  2, we propose a new paradigm for encoding circuit graphs, with a position-aware Hierarchical Transformer that enables operator-specific interaction among input variables, faithfully mirroring the computational process illustrated in Figure 2.
@@
A circuit graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$, consists of primary input (PI) nodes, which have an in-degree of zero, and operator nodes, whose in-degree is determined by their type. We begin by computing the logic level of each node in topological order [^1] To handle cyclic graphs, such as sequential AIGs, we follow Khan et al. (2025) by treating all flip-flops or registers as pseudo Primary Inputs (PIs) and removing the feedback edges to compute the logical level. as follows:
(2) $level(v)=\begin{cases}0&\text{if $v$ is PI}\\ 1+\max\limits_{(u,v)\in\mathbf{E}}level(u)&\text{otherwise}\end{cases}$
@@
The logic level defines the dataflow through the graph. Inspired by asynchronous MPNNs, we process nodes level by level. First, Primary Input (PI) nodes (level 0) are initialized with their input distribution, e.g., $\mathcal{B}(p)$, a Bernoulli distribution with parameter $p$. Then, for the nodes $v_{0}^{k},v_{1}^{k},\dots,v_{n_{k}}^{k}$ at subsequent level $k>0$, we gather their respective sets of direct predecessors, $\mathcal{N}(v_{0}^{k}),\dots,\mathcal{N}(v_{n_{k}}^{k})$, where each set is defined as $\mathcal{N}(v_{i}^{k})=\{u_{j}\in\mathbf{V}\mid(u_{j},v_{i}^{k})\in\mathbf{E}\}$.
@@
Inspired by prefix notation, we represent each computation at level $k$ as an ordered sequence, $\mathbf{v}_{i}^{k}$, formed by placing the operator $v_{i}^{k}$ at the head, followed by its ordered input nodes from $\mathcal{N}(v_{i}^{k})$:
(3) $\mathbf{v}_{i}^{k}=[v_{i}^{k},u_{j_{1}},u_{j_{2}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}],\text{ where }u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}\in\mathcal{N}(v_{i}^{k})$
@@
(4) $v_{i}^{k}=\text{Transformer}(\mathbf{v}_{i}^{k}+\mathbf{pos})[0],$
@@
This paradigm offers several key advantages. First, positional encodings enable position-aware aggregation for order-dependent operations, contrasting with the permutation-invariant nature of standard message passing. Second, self-attention enables rich interactions among all source nodes $u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}$], faithfully mirroring the computational dataflow of an operator. Finally, the sequence length depends only on in-degree: unlike general graphs with long-tailed degree distributions that incur high overhead, circuit graphs have small, tightly bounded in-degrees, preventing the memory explosion caused by quadratic complexity.
An Alternative View: Hierarchical Transformer on Prefix Notation. Our proposed method can also be interpreted as a Hierarchical Transformer operating on prefix notation, in a way that follows the logical dependencies inherent in the circuit graph. This perspective highlights how the model processes the graph by mirroring the step-by-step evaluation of an expression. A circuit graph can be converted into a prefix notation string through a pre-order traversal on its reversed edges. As illustrated in Figure 3, our proposed Hierarchical Transformer, unlike methods that simply flatten the graph into a single sequence, inherently preserves the nested, hierarchical structure of dependencies within each computation step, which allows it to align naturally with the actual flow of computation. This view also underscores the generalizability of our approach to other sequence-based problems, such as hardware model-checking, where the problem is often represented in Btor2 format (ArminBiere et al., 2018), a prefix-style representation of bit-vector formulas.
3.2. Function Shift Learning on Predictive Task
Logic-1 probability prediction is a task widely studied in prior works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2024; Zheng et al., 2025; Shi et al., 2025b; Liu et al., 2024), as it serves as a key indicator of a model’s ability to capture circuit functionality. The logic-1 probability corresponds to the global function of a circuit, as defined in Definition 3.1. This function can be highly complex due to reconvergent dependencies, and computing it directly requires enumerating the joint distribution of all inputs, which incurs an exponential cost of $O(2^{k})$.